OpenStudy (anonymous):

A pyramid has a height of 5 in and a surface area of 90 in2. Find the surface area for the similar pyramid with the height of 10 in. Round to nearest tenth of necessary @iGreen

OpenStudy (igreen):

Set up a proportion again. \(\dfrac{5}{90} = \dfrac{10}{x}\) Cross multiply. 5 * x = 90 * 10

OpenStudy (igreen):

@samsteronbroadway

OpenStudy (anonymous):

90 * 10 = 900

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yes, and 5 * x = 5x. So we have: 5x = 900 Divide 5 to both sides, what's 900 / 5?

OpenStudy (anonymous):

180

OpenStudy (igreen):

Yes, that's our answer.

OpenStudy (anonymous):

Thank you!